You need to use your cell phone, which broadcasts an 800 MHz signal, but you're behind two massive, radio-wave-absorbing buildings that have only a 15 m space between them. What is the angular width, in degrees, of the electromagnetic wave after it emerges?
Answer #1
Concepts and reason

The concepts needed to solve this problem are single slit diffraction and relation between the frequency, speed, and wavelength of a wave.

Initially, find the wavelength of the broadcasting signal using the frequency of the signal and speed of light. After that find, calculate the half angular width that is diffraction angle θ\theta by using diffraction minimum of single slit diffraction.

To find the angular width of the electromagnetic wave, multiply the angle θ\theta by a factor of two.

Fundamentals

The wavelength of the broadcasting signal is calculated as follows:

λ=cf\lambda = \frac{c}{f}

Here, c is the speed of light and f is the frequency of the broadcasting signal.

The expression for the diffraction minimum in the single slit diffraction is,

dsinθ=mλd\sin \theta = m\lambda

Here, d is the slit width, m is the order of the diffraction, andλ\lambda is the wavelength.

The angular width of the electromagnetic wave that is width of the central maximum is equal to the twice of the angle of diffraction for the first minimum.

θ=2θ\theta ' = 2\theta

The wavelength of the broadcasting signal is calculated as follows:

λ=cf\lambda = \frac{c}{f}

Substitute 800 MHz for f and 3.0×108m/s3.0 \times {10^8}{\rm{ m/s}}for c in the above expression.

λ=3.0×108m/s800MHz(106Hz1.0MHz)=0.375m\begin{array}{c}\\\lambda = \frac{{3.0 \times {{10}^8}{\rm{ m/s}}}}{{800{\rm{ MHz}}\left( {\frac{{{{10}^6}{\rm{ Hz}}}}{{1.0{\rm{ MHz}}}}} \right)}}\\\\ = 0.375{\rm{ m}}\\\end{array}

Rearrange the equationdsinθ=mλd\sin \theta = m\lambda forθ\theta .

θ=sin1(mλd)\theta = {\sin ^{ - 1}}\left( {\frac{{m\lambda }}{d}} \right)

The value of m for the first minimum is equal to one.

Substitute 1 for m, 0.375 m for λ\lambda and 15 m for d in the expressionθ=sin1(λd)\theta = {\sin ^{ - 1}}\left( {\frac{\lambda }{d}} \right).

θ=sin1((1)(0.375m)15m)=1.4\begin{array}{c}\\\theta = {\sin ^{ - 1}}\left( {\frac{{\left( 1 \right)\left( {{\rm{0}}{\rm{.375 m}}} \right)}}{{{\rm{ 15 m}}}}} \right)\\\\ = 1.4^\circ \\\end{array}

Substitute1.41.4^\circ forθ\theta in the equationθ=2θ\theta ' = 2\theta .

θ=2θ=2(1.4)=2.8\begin{array}{c}\\\theta ' = 2\theta \\\\ = 2\left( {1.4^\circ } \right)\\\\ = 2.8^\circ \\\end{array}

Ans:

The angular width of the electromagnetic wave is2.82.8^\circ .